In a division sum the dividend is 5045 and the divisor is 45 find the quotient?
Answer by Theo(13342)

the quotient is equal to the dividend divided by the divisor.

5045 / 45 = 112 with a remainder of 5

that's because 45 * 112 = 5040 and you have 5 left over.

you can also show it as 5045 / 45 = 112.1111.....

.11111 is equal to 5/45 which can be simplified to 1/9.

the quotient can therefore be shown as 112 and 1/9